A baseball is hit so that its height, h ( in feet), t seconds after it is hit can be represented by the equation h=-16t^2+56t+4
a) What is the height of the ball when it is hit ?

Answer by lwsshak3(11628)

A baseball is hit so that its height, h ( in feet), t seconds after it is hit can be represented by the equation h=-16t^2+56t+4
a) What is the height of the ball when it is hit ?
***
h=-16t^2+56t+4
complete the square:
h=-16(t^2-(56/16)t+4
h=-16(t^2-(7/2)t+49/16)+49+4
h=-16(t-7/4)^2+53
height of the ball when it is hit ?
53 ft after 1.75 sec